OpenStudy (anonymous):

A girl asks her father how many ties he has and he answers with a riddle: "Seven-eighths of my ties plus five." How many ties does her father have?

OpenStudy (animallover8):

This doesn't make sense. I think you left something out.

OpenStudy (mathstudent55):

He has x ties. He says he has 7/8 x + 5 ties That means x = 7/8 x + 5 Solve the equation for x.

OpenStudy (mathstudent55):

x = 7/8 x + 5 Subtract 7/8 x from both sides: x - 7/8 x = 5 8/8 x - 7/8 x = 5 1/8 x = 5 x = 40